Advances in thrill ride technologies have made electromagnetic slingshots possible instead of using large hills to propel roller coasters. Suppose a 500kg roller coaster is accelerated by a 200 kW electromagnetic motor for 4 seconds, what is its final velocity?
1
Expert's answer
2019-05-06T10:20:28-0400
"E=Pt=\frac{mv^2}{2}"

"(200000)(4)=500\frac{v^2}{2}"

"v=56.6\frac{m}{s}"